OpenStudy (anonymous):

12a^2-13a=35 solve the equation check your solution show work

OpenStudy (anonymous):

a = -5/4 a = 7/3

OpenStudy (anonymous):

12a² - 13a = 35 12a² - 13a - 35 = 0 12a² - 28a + 15a - 35 = 0 4a(3a - 7) + 5(3a - 7) = 0 (4a+5)(3a-7) = 0 a = -5/4 or a = 7/3

OpenStudy (anonymous):

12a^2 - 13a = 35 so subtract 35 from both sides: 12a^2 - 13a - 35 = 0 Factor: (4a + 5)(3a - 7) a = -5/4 a = 7/3